# Improving Generalization via Meta-Learning on Hard Samples

### v1

@@ 1 Introduction
  • We show strict accuracy ordering of LRW classifiers based on validation set: easy $<$ random $<$ hard, clearly demonstrating the importance of optimizing LRW validation sets. We obtain reliable gains over ERM across datasets (e.g., 1% on Imagenet/VIT-B backbone), and in domain-generalization (e.g., 1.36% on iWildCam dataset [4] from Wilds benchmark [19]) and noisy-label settings (e.g., 4.2% on Clothing1M dataset [40]). We outperform a range of baselines including reweighting, meta-learning among others. Analyses show improved margins on test set in MOLERE classifiers compared to ERM, suggesting a mechanism by which MOLERE provides gains.
  • We extend these results to natural hard samples (Imagenet training with Imagenet-R / Imagenet-A) as validation, showing 1-3% gains on both in-domain and out-of-domain test sets (Imagenet, and Imagenet-A/Imagenet-R respectively). This shows the value of our ideas even in scenarios where we pay no additional cost for meta-optimization.
@@ 2 Related Work
Other than these loss-based importance methods, a recent line of work [17, 18] proposed weighting based on context/relevance of an instance compared with the overall data distribution. These papers target slow temporal drift in longitudinal datasets, and the development of better uncertainty measures for selective classification, respectively. In particular, Jain et al. [18], suggest that using target domain data in the validation set can improve domain shift performance of classifiers.
Meta Learning. The sample re-weighting task using a validation set is comes under the umbrella of meta-learning [16], which follows the learning-to-learn paradigm. It is highly similar to a model agnostic meta learning (MAML) [11] setup, which involves updating the parameters using the multi-task setup such that performance of the parameters updated according to a particular task is maximized on the test set for that task, leading to two optimization loops idealistically. On similar lines, are these re-weighting techniques which aim to update the meta network based on classifier performance of this meta-test set. Lately, there have been significant developments [32, 30, 45] in order to make this optimization process more efficient.
Probabilistic Margins. A recent work [24] introduced probabilistic margins of neural networks in multi-class setups and showed how it can be an efficient geometric measurement when dealing with adversarial examples. These margins in some sense account for the toughness of a given instance and are basically the difference in predicted probabilities of the true label and the highest probability class excluding the true label class. The method involved re-weighting training instances, in presence of an adversarial attack, inversely related to their probability margins.
@@ 4.2 MOLERE objective and generalization
Consider the tri-level optimization in Equation (2). Suppose the weighting function ϕ(),italic-$\phi(\cdot),$ and splitting function Θ()normal-$\Theta(\cdot)$ are dependent on both $x$ and $y$. Let’s suppose $N+M\to\infty$, and $\lim_{N,M\to\infty}\frac{M}{N+M}=\delta$. Moreover, suppose the domains of $\phi,\theta,\Theta$ are very large and contain the set of all measurable functions. Then the objective of MOLERE  is equivalent to
@@
The proof of the theorem relies on the observation that there exists a weighting function $\phi$ that can transform any probability distribution $P$ to any another distribution $Q$ (as long as the support of $Q$ is a subset of $P$). Using this observation, one can show that the second, third level optimization problems in Equation (2) is equivalent to following problem: $\min_{\theta}\sum_{(x,y)\in\Theta(S)^{c}}\ell(y,f_{\theta}(x))$. ∎
@@ 5.1 Datasets
We use popular classification benchmarks including CIFAR-100 [21], ImageNet-100 [37], ImageNet-1K [8], Aircraft [25], Stanford Cars [20], Oxford-IIIT Fine-grained classification (Cats v/s Dogs) [31] and Diabetic Retinopathy (DR) dataset. Furthermore, for OOD analysis we use the ImageNet-A [14], ImageNet-R [13] datasets for models trained on ImageNet-1K. We also use the Camelyon [2], iWildCam [4] dataset from the widely popular WILDS benchmark [19] and also the country shifted test set for the DR [1] dataset. We further analyze for robustness in presence of instance dependent noise on the noisy version of CIFAR-100 dataset (Inst.C-100) proposed by Xia et al. [38] and Clothing-1M [40] dataset. Please refer supplementary (supp.) for more details, setup for these datasets.
@@ 6.1.1 MOLERE improves classification accuracy
  • The end-to-end optimization method LRWOpt mostly matches or exceeds LRW-Hard, showcasing its effectiveness without the need for training twice. On a minority of datasets, LRW-Hard is nominally better; we believe this is due to the small-sample nature of real-world datasets, and attendant estimation noise.
@@ 6.2 Out-Of-Distribution (OOD) generalization
6.3 Practical Noisy Scenarios
We now evaluate our method for a noisy label case which is a usual problem in classification datasets. However keeping the real-world applications in mind, we use datasets like Clothing1M [40], noisy CIFAR-100 proposed in [38] (Inst. C-100) which contain instance conditioned noise as against randomly flipping the labels or adding uniform noise. This noise can be treated as a measure of hardness to classify an instance due to its characterstics. Please refer supplementary for more details on these datasets. Table 3 provides the analysis of our LRWOpt method against the bi-level optimization based baselines like MWN, FSR, L2R, GDW specifically designed for noisy label case but usually tested on label flipping, uniform noise cases. We also compare against the MAPLE baseline since it is designed for general robustness. On the Inst. C-100 dataset, our method clearly surpasses all of these methods, showing 1.53% gain in accuracy. On the Clothing-1M, our method shows about 0.85% gain over the best performing GDW baseline. For other noisy settings like random label flips or high degree of noise, prior work on outlier robust DRO [41] has shown good results on noisy data using loss clipping (e.g. treat kof highest loss as noisy instance to be excluded). We could rely on a similar techniques to handle corruptions in our framework.
6.4 Skewed Label Setup
We now analyze our LRWOpt proposition on a skewed label scenario against existing instance based reweighting schemes (Meta-Weight-Net, FSR and another recently propoed re-weighting method GDW [6]) which have been proposed with this setting as a key use-case. Table 2 shows the results for this analysis on the CIFAR-100 datasets with various skew levels ranging from 1 to 200. It can be observed that our LRWOpt significantly outperform the existing loss based reweighting methods at all of the skew levels showing accuracy gains upto 2.22%.
@@ 6.5 MOLERE scales to large pretrained models
Figure 3: MOLERE improves margins of learned classifiers. (a,b): paired margin deltas between LRWOpt and ERM are moderately right-skewed with mean/median greater than zero. (c,d): As a function of ERM margin, clear separation seen between LRW-Hard (better) and LRW-Easy (worse) in terms of margin gain over ERM (errorbars are SEM). All results on unseen test data; Imagenet-100, CIFAR100 shown for brevity, with similar results for other datasets in appendix.
@@ 6.7 Margin maximization via meta-learning
We present empirical evidence showing that MOLERE has a margin maximization effect, i.e., the learned classifiers have wider margins than ERM classifiers. Since our validation data is selected to be low-margin instances, our LRW classifier upweights, and improves performance on, training instances most similar to the low-margin validation set. 3 shows two views of margin differences between MOLERE and ERM on the test set, to confirm this expectation [^3] We show Imagenet-100 & CIFAR100 for brevity; findings consistent across all datasets (see Appendix).. Panels (a,b) shows a histogram of paired margin differences between ERM and LRWOpt, indicating a modest right-skew with mean & median to the right of 0. Panels (c,d) shows LRW-Hard and LRW-Easy deltas w.r.t. ERM, averaged over ERM margin buckets (mean and S.E.M. errorbars)–across the board (i.e., for most values of ERM margin), LRW-Hard contrasts with ERM better than LRW-Easy.
@@ Appendix B Proof of Theorem $1$
Consider the tri-level optimization in Equation $(2)$. Suppose the weighting function ϕ(),italic-$\phi(\cdot),$ and splitting function Θ()normal-$\Theta(\cdot)$ are dependent on both $x$ and $y$. Let’s suppose $N+M\to\infty$, and $\lim_{N,M\to\infty}\frac{M}{N+M}=\delta$. Moreover, suppose the domains of $\phi,\theta,\Theta$ are very large and contain the set of all measurable functions. Then the objective of MOLERE  is equivalent to
@@
Observe that the above two optimization problems are the inner optimization problems of objectives (2) and (8). Showing that these two are equivalent would then immediately imply that objective (2) is equivalent to objective (8).
First let’s consider the case where $\text{supp}(Q^{\text{tr}})=\text{supp}(Q^{\text{val}})$. By choosing $\phi(x,y)=\frac{Q^{\text{val}}(x,y)}{Q^{\text{tr}}(x,y)}$, the constraint in Equation (10) can be rewritten as
@@
Observe that this is the same as the optimization problem in Equation (9). This shows that both the optimization problems in Equation (9), (10) are equivalent.
Next, consider the case where $\text{supp}(Q^{\text{tr}})\neq\text{supp}(Q^{\text{val}})$. This is the easy case to handle. To see this, consider the extreme case where $\text{supp}(Q^{\text{tr}})\cap\text{supp}(Q^{\text{val}})=\{\}$. Since the domain of $\theta$ contains the set of all measurable functions, we can set $\phi(x,y)=1$ and choose $\theta^{*}(\phi)$ to be the Bayes optimal classifier [^4] A Bayes optimal classifier is a classifier that minimizes the expected population risk on both $\text{supp}(Q^{\text{tr}}),\text{supp}(Q^{\text{val}})$. It is easy to verify that this is an optimizer of Equation (10). This shows that Equation (9), (10) are equivalent. A similar argument can be used to handle the more general case of $\text{supp}(Q^{\text{tr}})\neq\text{supp}(Q^{\text{val}})$. Here, we choose a $\phi$ that performs probability matching on the intersection of the two supports, and set $\phi(x,y)=1$ on the rest of the support. ∎
@@ D.2 Datasets
Stanford Cars [20]. This dataset contains around 16k images from 102 categories of cars based on Model, company, etc. with a roughly equal split between train and test set. We use 1K images (around 10 images per class), from the train set use as our validation set.
@@ Appendix F Comparison with only hard examples in validation set
We further analyze three new variants which involve using only hard examples in the validation set: First we incorporate the loss highlighted in [22] for the validation set along with our LRW-Hard method. Second, in our LRWOpt method we decrease the threshold $\Theta$ for train set to 0.2, such that only the hardest of the examples are there in the validation set and third where in the LRW-hard, we limit the validation set to only negative margin (i.e., incorrectly classified) examples from ERM. Table 7 below shows accuracy % gains of LRWOpt over these variants on 4 randomly picked datasets including 1 OOD challenge.

### v2

@@ 1 Introduction
  • We show strict accuracy ordering of LRW models based on validation set: easy $<$ random $<$ hard, demonstrating the importance of optimizing LRW validation sets. We obtain reliable gains over ERM across datasets (e.g., 1% on Imagenet/VIT-B backbone), and in domain-generalization (e.g., 1.36% on iWildCam dataset [4, 19]) and noisy-label settings (e.g., 4.2% on Clothing1M dataset [40]). We outperform a range of baselines including reweighting, meta-learning among others. Analyses show improved margins on test set in MOLERE classifiers, suggesting an explanation of our gains.
  • We extend our results to natural hard samples as validation (Imagenet with Imagenet-R / Imagenet-A), showing 1-3% gains on both in-domain and out-of-domain test sets. This shows the value of our ideas even in scenarios where we pay no additional cost for meta-optimization.
@@ 2 Related Work
A recent line of work [17, 18] proposed weighting based on context/relevance of an instance compared with the overall data distribution. These papers target slow temporal drift in longitudinal datasets, and the development of better uncertainty measures for selective classification, respectively. In particular, Jain et al. [18], suggest that using target domain data in the validation set can improve domain shift performance of classifiers.
Meta Learning. The sample re-weighting task using a validation set comes under the umbrella of meta-learning [16], which follows the learning-to-learn paradigm. It is conceptually similar to model agnostic meta learning (MAML) [11], which learns a single set of parameters that can easily be customized (few-shot) to multiple tasks. Learning these shared parameters involves a nested optimization similar to the one presented here, and significantly optimized for efficiency by recent work[32, 30, 45].
Probabilistic Margins. Recent work [24] showed that the probabilistic margin in multi-class problems can be used to improve how neural networks deal with adversarial examples. The margin is defined as the difference between the probability of the true label and the largest of the remaining label probabilities, indicating the difficulty of classifying that instance. Liu et al. [24] propose re-weighting training instances, in presence of an adversarial attack, inversely related to their probability margins.
@@ 4.2 MOLERE objective and generalization
Consider the tri-level optimization in Equation (LABEL:eqn:tri_level). Suppose the weighting function $\phi(\cdot),$ and splitting function $\Theta(\cdot)$ are dependent on both $x$ and $y$. Let’s suppose $N+M\to\infty$, and $\lim_{N,M\to\infty}\frac{M}{N+M}=\delta$. Moreover, suppose the domains of $\phi,\theta,\Theta$ are very large and contain the set of all measurable functions. Then the objective of MOLERE  is equivalent to
@@
The proof of the theorem relies on the observation that there exists a weighting function $\phi$ that can transform any probability distribution $P$ to any another distribution $Q$ (as long as the support of $Q$ is a subset of $P$). Using this observation, one can show that the second, third level optimization problems in Equation (LABEL:eqn:tri_level) is equivalent to following problem: $\min_{\theta}\sum_{(x,y)\in\Theta(S)^{c}}\ell(y,f_{\theta}(x))$. ∎
@@ 5.1 Datasets
We use popular classification benchmarks including CIFAR-100 [21], ImageNet-100 [37], ImageNet-1K [8], Aircraft [25], Stanford Cars [20], Oxford-IIIT Fine-grained classification (Cats v/s Dogs) [31] and Diabetic Retinopathy (DR) dataset. Furthermore, for OOD analysis we use the ImageNet-A [14], ImageNet-R [13] datasets for models trained on ImageNet-1K. We also use the Camelyon [2], iWildCam [4] dataset from the widely popular WILDS benchmark [19] and also the country shifted test set for the DR [1] dataset. We further analyze for robustness in presence of instance dependent noise on the noisy version of CIFAR-10 dataset (Inst.C-10) proposed by Xia et al. [38] and Clothing-1M [40] dataset. Please refer supplementary (supp.) for more details, setup for these datasets.
@@ 6.1.1 MOLERE improves classification accuracy
  • The end-to-end LRWOpt matches or exceeds LRW-Hard, showcasing its effectiveness without the need for training twice. On a minority of datasets, LRW-Hard is nominally better; we believe this is due to the small-sample nature of real-world datasets, and attendant estimation noise.
@@ 6.2 Out-Of-Distribution (OOD) generalization
6.3 Practical Label Noise Settings
We now study the noisy label setting with a focus on real-world label noise using datasets like Clothing1M [40], and noisy CIFAR-10 (Inst. C-10 [38]), which contain instance conditioned noise as against randomly flipping labels or adding uniform noise. This noise can be treated as a measure of instance hardness rather than labeling error (refer supplementary for more details). Table 3 compares LRWOpt method against bi-level optimization approaches like MWN, FSR, L2R, GDW which although specifically designed for noisy labels have been primarily tested on label flipping, uniform noise data. We also compare against MAPLE, designed for general robustness. On Inst. C-10, our method clearly surpasses all baselines, with 1.53% accuracy gain. On Clothing-1M, we show about 0.85% gain over the best performing GDW baseline. For other noisy settings like random label flips, prior work on outlier robust DRO [41] has shown good results using loss clipping (e.g. exclude kof highest loss as label noise), which could be co-opted into our work.
6.4 Skewed Labels
We now study the skewed label setting, comparing against instance based reweighting schemes proposed with this setting in mind (Meta-Weight-Net, FSR and another recently proposed re-weighting method GDW [6]). Table 2 shows the results for this analysis on the CIFAR-100 datasets with various skew levels ranging from 1 to 200. It can be observed that our LRWOpt significantly outperform the existing loss based reweighting methods at all of the skew levels showing accuracy gains upto 2.22%.
@@ 6.5 MOLERE scales to large pretrained models
Figure 3: MOLERE improves margins of learned classifiers. (a,b): paired margin deltas between LRWOpt and ERM are moderately right-skewed with mean/median greater than zero. (c,d): As a function of ERM margin, clear separation seen between LRW-Hard (better) and LRW-Easy (worse) in terms of margin gain over ERM (errorbars are SEM). All results on unseen test data; Imagenet-100, CIFAR100 shown for brevity, with similar results for other datasets in supplementary.
@@ 6.7 Margin maximization via meta-learning
We present empirical evidence showing that MOLERE has a margin maximization effect, i.e., the learned classifiers have wider margins than ERM classifiers. Since our validation data is selected to be low-margin instances, our LRW classifier upweights, and improves performance on, training instances most similar to the low-margin validation set. 3 shows two views of margin differences between MOLERE and ERM on the test set, to confirm this expectation [^3] We show Imagenet-100 & CIFAR100 for brevity; findings consistent across all datasets (see supplementary).. Panels (a,b) shows a histogram of paired margin differences between ERM and LRWOpt, indicating a modest right-skew with mean & median to the right of 0. Panels (c,d) shows LRW-Hard and LRW-Easy deltas w.r.t. ERM, averaged over ERM margin buckets (mean and S.E.M. errorbars)–across the board (i.e., for most values of ERM margin), LRW-Hard contrasts with ERM better than LRW-Easy.
Further details on Algorithm 1, the time complexity of the proposed scheme compared to ERM, the training setup, ablation study over the validation set selection/loss, and discussion regarding the early-stage performance of the proposed scheme are provided in the supplementary.
@@ Appendix B Proof of Theorem $1$
Consider the tri-level optimization in Equation $(2)$. Suppose the weighting function $\phi(\cdot),$ and splitting function $\Theta(\cdot)$ are dependent on both $x$ and $y$. Let’s suppose $N+M\to\infty$, and $\lim_{N,M\to\infty}\frac{M}{N+M}=\delta$. Moreover, suppose the domains of $\phi,\theta,\Theta$ are very large and contain the set of all measurable functions. Then the objective of MOLERE  is equivalent to
@@
Observe that the above two optimization problems are the inner optimization problems of objectives (LABEL:eqn:tri_level) and (8). Showing that these two are equivalent would then immediately imply that objective (LABEL:eqn:tri_level) is equivalent to objective (8).
First let’s consider the case where $\text{supp}(Q^{\text{tr}})=\text{supp}(Q^{\text{val}})$. By choosing $\phi(x,y)=\frac{Q^{\text{val}}(x,y)}{Q^{\text{tr}}(x,y)}$, the constraint in Equation (LABEL:eqn:trilevel_core) can be rewritten as
@@
Observe that this is the same as the optimization problem in Equation (9). This shows that both the optimization problems in Equation (9), (LABEL:eqn:trilevel_core) are equivalent.
Next, consider the case where $\text{supp}(Q^{\text{tr}})\neq\text{supp}(Q^{\text{val}})$. This is the easy case to handle. To see this, consider the extreme case where $\text{supp}(Q^{\text{tr}})\cap\text{supp}(Q^{\text{val}})=\{\}$. Since the domain of $\theta$ contains the set of all measurable functions, we can set $\phi(x,y)=1$ and choose $\theta^{*}(\phi)$ to be the Bayes optimal classifier [^4] A Bayes optimal classifier is a classifier that minimizes the expected population risk on both $\text{supp}(Q^{\text{tr}}),\text{supp}(Q^{\text{val}})$. It is easy to verify that this is an optimizer of Equation (LABEL:eqn:trilevel_core). This shows that Equation (9), (LABEL:eqn:trilevel_core) are equivalent. A similar argument can be used to handle the more general case of $\text{supp}(Q^{\text{tr}})\neq\text{supp}(Q^{\text{val}})$. Here, we choose a $\phi$ that performs probability matching on the intersection of the two supports, and set $\phi(x,y)=1$ on the rest of the support. ∎
@@ D.2 Datasets
Inst. C-10[38]. The setup here is same as the CIFAR-100 dataset with 30% noise. Stanford Cars [20]. This dataset contains around 16k images from 102 categories of cars based on Model, company, etc. with a roughly equal split between train and test set. We use 1K images (around 10 images per class), from the train set use as our validation set.
@@ Appendix F Comparison with only hard examples in validation set
We further analyze three new variants which involve using only hard examples in the validation set: First we incorporate the loss highlighted in [22] for the validation set along with our LRW-Hard method. Second, in our LRWOpt method we decrease the threshold $\Theta$ for train set to 0.2, such that only the hardest of the examples are there in the validation set and third where in the LRW-hard, we limit the validation set to only negative margin (i.e., incorrectly classified) examples from ERM. Table 7 shows accuracy % gains of LRWOpt over these variants on 4 randomly picked datasets including 1 OOD challenge.